OpenStudy (anonymous):

Solve: −5(x − 2) = 15 1 −3.4 −5 −1

OpenStudy (anonymous):

−5(x − 2) = 15 DISTRIBUTE THE -5 TO EVERYTHING INSIDE THE PARENTHESIS

OpenStudy (anonymous):

c

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

-5(x-2) = 15 -5 * x = -5x -5 * -2 = +10 So: -5x + 10 = 15 SUBTRACT THE 10 FROM BOTH SIDES -5x + 10 = 15 -10 -10 ----- ----- -5x = 5 DIVIDE BY -5 ON BOTH SIDES -5x = 5 --- --- -5 -5 x = ____

OpenStudy (anonymous):

@_litoswave_ do you understand? what did you get now? :)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

negative 1 my fault

OpenStudy (anonymous):

YES :)